OpenStudy (anonymous):

1. How long does money in a savings account take to double itself at 4% interest compounded annually? [use the formula A=P(1+i)^n]. Round answer to nearest tenth.

OpenStudy (anonymous):

You need to find n so that \[ P(1+i)^ n = 2 P\\ 1.04^n= 2\\ \] We cancelled by P. Can you finish it?

OpenStudy (anonymous):

Im lost. i got 50.99346 and my answer choices are 1.9 3.1 5.7 17.7 23.2

OpenStudy (anonymous):

Sorry, this is the correct solution for n \[ n \ln (1.04) =\ln(2)\\ n = \frac {\ln(2)}{\ln (1.04)} \]

OpenStudy (anonymous):

\[n =\frac {0.693147}{0.0392207}=17.673 \] The answer 17.7